OpenStudy (shamil98):

The greatest integer, [x], is equal to the largest integer that is smaller than or equal to x. In the domain [0,100], how many solutions are there to: \[x^2 - [x]x = 81.25\]

OpenStudy (the_fizicx99):

the answer is D :)

OpenStudy (shamil98):

fizi stop trolling -.-

OpenStudy (anonymous):

100-81.25

OpenStudy (shamil98):

why though batman

OpenStudy (the_fizicx99):

k :/

OpenStudy (anonymous):

k thanks bye close

OpenStudy (shamil98):

dumb it down for me and explain

OpenStudy (anonymous):

Well, factoring might help: \[ x(x-\lfloor x\rfloor)=81.25 \]

OpenStudy (anonymous):

legit way ^

OpenStudy (anonymous):

Suppose \(x = \lfloor x\rfloor +a\). We know \(0\leq a < 1\)

OpenStudy (anonymous):

Also:\[ x-\lfloor x\rfloor = a \]

OpenStudy (anonymous):

so \[ xa = 81.25 \]

OpenStudy (anonymous):

\[ a = \frac{81.25}{x}\implies 0\leq \frac{81.25}{x}<1 \]

OpenStudy (anonymous):

So \[ 81.25 < x\leq 100 \]

OpenStudy (shamil98):

ah, makes sense, ty.

OpenStudy (anonymous):

Sham get off that website, if you're going to cheat.

OpenStudy (anonymous):

<3

OpenStudy (shamil98):

this is about learning batman shutup

OpenStudy (shamil98):

if i was aiming for just an answer i would'nt have asked for an explanation

OpenStudy (anonymous):

I was joking

OpenStudy (anonymous):

I wasn't sure what I was doing. I winged it.

OpenStudy (anonymous):

It's pretty much right lol

OpenStudy (shamil98):

worked out in the end xD

OpenStudy (anonymous):

Did you check it?

OpenStudy (anonymous):

Your answer should be 19 sham, ye?

OpenStudy (anonymous):

My solution is not complete.

OpenStudy (anonymous):

I only really gave a feasible range, but there will be many gaps in it.

OpenStudy (anonymous):

I didn't put restrictions on \(a\) despite the fact that \(a=x-\lfloor x\rfloor\)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

if we said \(\lfloor x\rfloor =82\) then:\[ a = \frac{81.25}{82} \]

OpenStudy (shamil98):

yeah, they wanted the integer solution so i put 19 batman.

OpenStudy (anonymous):

So \(x =82+ 81.25/82\) is perhaps a solution... the reasoning make sense but it's important to verify.

OpenStudy (anonymous):

What? If \(x\) is an integer then \(x^2-x\lfloor x \rfloor = x^2-x^2=0\)

OpenStudy (anonymous):

I would say all solutions are: \[ \left\{\lfloor x\rfloor +\frac{81.25}{\lfloor x\rfloor }\bigg|x\in (81.25,100]\right\} \]

OpenStudy (anonymous):

Are you guys talking about a different problem?

OpenStudy (anonymous):

Same

OpenStudy (anonymous):

Why 19?

OpenStudy (anonymous):

One small alteration:\[ \left\{\lfloor x\rfloor +\frac{81.25}{ x }\bigg|x\in (81.25,100]\right\} \]

OpenStudy (anonymous):

x^2-[x]x=81.25 x-[x]=81.25/x

OpenStudy (anonymous):

Why 19 though?